# Momentum coupling of classical atoms

### v2

@@ 3 Thermodynamic limit
Infinite clusters only appear in this case, for $\rho>\rho_{c}(\beta)$, signaling that finite clusters cannot carry a density greater than $\rho_{c}(\beta)$. The argument we gave above applies to interactions without a hard core. Now $u$ with a hard core is allowed if we apply a mollifier to the Boltzmann factor $e^{-\beta u(x)}$. In this case $\rho_{c}(\beta)\leq\rho_{\rm cp}$, the close-packing density, and the condition for the appearance of infinite clusters is $\rho_{c}(\beta)=\bar{A}_{\beta}/\lambda_{\beta}^{d}<\rho_{\rm cp}$. $A_{\rho,\beta}(\mu)$ cannot be continued analytically beyond $\rho_{\rm cp}\lambda_{\beta}^{d}$, so the alternative is $\bar{A}_{\beta}/\lambda_{\beta}^{d}=\rho_{\rm cp}$, no infinite cluster for $\rho<\rho_{\rm cp}$. Note in parentheses that if $\rho<\rho_{c}(\beta)$, $e^{-n\rho\hat{v}(0)}q_{n}e^{n\beta\partial f/\partial\rho}$ decays at least exponentially fast, and from Eq. (10)
$\frac{\partial^{2}f(\rho,\beta)}{\partial\rho^{2}}=\left(\frac{\partial\rho}{\partial\mu}\right)^{-1}_{\mu=\partial f/\partial\rho}=\frac{\lambda_{\beta}^{d}}{\partial A_{\rho,\beta}(\mu)/\partial\mu\left|{}_{\mu=\partial f/\partial\rho}\right.}=\frac{\lambda_{\beta}^{d}}{\beta\sum_{n=1}^{\infty}ne^{-n\rho\hat{v}(0)}q_{n}e^{n\beta\partial f(\rho,\beta)/\partial\rho}}>0,$ (15)
due to the convergence of the infinite sum. The interest in $\partial^{2}f(\rho,\beta)/\partial\rho^{2}$ lies in its relation with the $\rho$-derivative of the pressure:
@@
We now derive the $L\to\infty$ limit of $q^{L}_{n}$. Clearly, $q_{1}=1$ and $q_{2}=0$. For $n\geq 3$ we proceed as follows. Substituting $E_{L}(x)=1-v_{L}(x)$ in the formula for $\hat{E}_{L}(z)$,
@@
The dependence on $L^{d}$ shows that among $\{z^{j}_{i}\in{\mathbb{Z}}^{d}\setminus\{0\}|(ij)\in E(C)\}$ there are $N_{\rm free}(C)=|E(C)|-n+1$ free summation variables, and the residual $n-1$ variables are linear combinations of them. In the limit of $L$ going to infinity, $L^{-d(|E(C)|-n+1)}$ times the multiple sum over $\{z^{j}_{i}\in{\mathbb{Z}}^{d}\setminus\{0\}|(ij)\in E(C)\}$ becomes a multiple integral with respect to the free variables,
@@
This is in accordance with
@@
Remark 1. If a block is a planar graph or a graph consisting of the vertices and edges of a convex polyhedron, $N_{\rm free}$ is related to the Euler characteristic $V+F-E=2$. Comparison with it shows that $N_{\rm free}=F-1$, where $F$ is the number of faces (the exterior face included for planar graphs). We choose any $F-1$ faces of the polyhedron, or all the interior faces of the planar block, and assign a free variable to each bordering cycle. Referring to Eq. (17), $c_{ij}\neq 0$ for a single $j$ if $i$ is an external edge and for two $j$’s if $i$ is an internal edge.
Remark 2. For the complete $n$-graph the general solution of the system of equations $Z_{k}=0$, $k=1,\dots,n$, is easily obtained. In this case $N_{\rm free}=n(n-1)/2-n+1=(n-1)(n-2)/2$, the number of edges of the complete $(n-1)$-graph. The variables of this latter, $z^{k}_{j}$ for $11\leq j<k\leq n-$ 1 can be taken to be free; then, with
@@
all the equations will be satisfied. The integration variables are $\kappa^{k}_{j}$, $11\leq j<k\leq n-$ 1.
@@ 4 An illustrative example
$u$ depends on $\beta$, but as long as we only look at the $\rho$-dependence, the interaction is legitimate. Let us retain from $\Sigma_{n}$ only the $(n-1)!/2$ cycles of length $n$. There are many more elements of $\Sigma_{n}$, and one can only hope that, with the cycles only, the result is qualitatively correct. There is no phase transition in one dimension for short-range repulsive interactions, so we focus on $d\geq 2$. For $n\geq 3$
@@
If $\beta\mu<d\ln(\lambda_{\beta}/\lambda)+\rho\hat{v}(0)$, the infinite sum is absolutely convergent. If $\beta\mu=d\ln(\lambda_{\beta}/\lambda)+\rho\hat{v}(0)$, it is still absolutely convergent for $d\geq 3$, nonabsolutely convergent for $d=1,2$, and for any $d\geq 1$ it is bounded below by $-3^{-d/2}$, so
@@
is convergent, by Abel’s theorem $A_{\rho,\beta}(\mu)$ tends to it as $\mu$ tends to $\bar{\mu}(\rho,\beta)$. It remains for us to verify the consistency of the approximation. First, $A_{\rho,\beta}(\mu)$ must monotonically increase with $\mu$ for $-\infty<\mu<\bar{\mu}(\rho,\beta)$; second, $\lim_{\mu\uparrow\bar{\mu}(\rho,\beta)}{\,\rm d}A_{\rho,\beta}(\mu)/{\,\rm d}\mu$ must exist, so that $\partial p(\rho,\beta)/\partial\rho$ has a left limit at $\rho_{c}(\beta)$, cf. Eqs. (15), (16). By introducing $\gamma=-\beta[\mu-\bar{\mu}(\rho,\beta)]$, $A_{\rho,\beta}(\mu)=(\lambda_{\beta}/\lambda)^{d}B(\gamma)$, where
@@ 5 No condensed phase without infinite clusters
Not knowing $q_{n}$ any better, we can nevertheless affirm that any condensed phase must contain infinite clusters. Consider, for example, the vapor-liquid transition. Fix a subcritical temperature and increase the density of the vapor by increasing the pressure. At some $\rho=\rho_{v+}(\beta)$ condensation begins and, as we know from experiment, the van der Waals theory, and numerical simulations with, for example, the Lennard-Jones potential [15, 16], the vapor of density $\rho_{v+}(\beta)$ and the liquid of density $\rho_{l-}(\beta)$ coexist for the global density in the interval $(\rho_{v+}(\beta),\rho_{l-}(\beta))$. The chemical potential is the same in both phases and is therefore constant throughout the coexistence interval. It is $\mu(\rho_{v+}(\beta),\beta)=\epsilon(\beta)+\rho_{v+}(\beta)\hat{v}(0)/\beta$, cf. Eqs. (11), (13), where
@@
is valid for any $\rho\geq\rho_{v+}(\beta)$. Beyond $\rho_{l-}(\beta)$ the chemical potential continues to increase with $\rho$, and the equation (25) defines a condition on it. By further increasing the pressure, we arrive at the coexistence of the liquid of density $\rho_{l+}(\beta)$ and a solid of density $\rho_{s-}(\beta)$. In the coexistence interval $(\rho_{l+}(\beta),\rho_{s-}(\beta))$ the chemical potential is again constant at the value $\partial f(\rho,\beta)/\partial\rho|_{\rho=\rho_{l+}(\beta)}$, and continues to increase beyond $\rho=\rho_{s-}(\beta)$. The infinite clusters in a solid must be different from those in the liquid. In a crystal the atomic positions are correlated on long distances, in the liquid the correlation is of medium range. This suggests that in a crystal the infinite clusters are macroscopic, while in the liquid they are submacroscopic. Accordingly, for $\rho\geq\rho_{v+}(\beta)$
@@
densities carried by finite, infinite submacroscopic, and macroscopic clusters. Submacroscopic clusters must be infinite in number to contribute to the density. If $\beta<\infty$ then $\rho_{v+}(\beta)>0$, finite clusters appear along with infinite clusters, in liquids and crystals, at positive temperatures. On the contrary, $\lim_{\beta\to\infty}\rho_{v+}(\beta)$ can be positive or zero, and both results have an interesting consequence on the ground state, the $\beta\to\infty$ limit of the infinite volume Gibbs state. As $\beta$ tends to infinity, $q_{n}$ tends to zero for $n\geq 3$, see the discussion following Eq. (20). Therefore, the convergence radius of the fugacity expansion, the first term in the right member of Eq. (14), diverges, but its increase must be compared to that of $\lambda_{\beta}^{d}$. If $\rho_{v+}(\infty)>0$, there are finite clusters in the ground state at any $\rho<\rho_{\rm cp}\leq\infty$, a possible result for short-range interactions. On the other hand, if $\rho_{v+}(\infty)=0$, there are only infinite clusters in the ground state at arbitrarily low density, a possible result for slowly decaying, but still integrable, interactions.

### v3

@@ 3 Thermodynamic limit
The above argument applies to interactions without a hard core. Now $u$ with a hard core is allowed if we apply a mollifier to the Boltzmann factor $e^{-\beta u(x)}$. In this case $\rho_{c}(\beta)\leq\rho_{\rm cp}$, the close-packing density, and the condition for the appearance of infinite clusters is $\bar{A}_{\beta}/\lambda_{\beta}^{d}<\rho_{\rm cp}$. $A_{\rho,\beta}(\mu)$ cannot be continued analytically beyond $\rho_{\rm cp}\lambda_{\beta}^{d}$, so the alternative is $\bar{A}_{\beta}/\lambda_{\beta}^{d}=\rho_{\rm cp}$, no infinite cluster for $\rho<\rho_{\rm cp}$.
Note that for $\rho<\rho_{c}(\beta)$, $e^{-n\rho\hat{v}(0)}q_{n}e^{n\beta\partial f/\partial\rho}$ decays at least exponentially fast, $f(\rho,\beta)$ is a real analytic function of $\rho$, and from Eq. (12) and the implicit function theorem a relation between the first and second $\rho$-derivative of $f(\rho,\beta)$ can be derived. Let
$F(\rho,\mu):=\rho\lambda_{\beta}^{d}-A_{\rho,\beta}(\mu),$
Differentiating $F(\rho,\mu(\rho))=0$ with respect to $\rho$ gives
$\frac{d\mu}{d\rho}=\frac{\lambda_{\beta}^{d}-\partial A_{\rho,\beta}(\mu)/\partial\rho}{\partial A_{\rho,\beta}(\mu)/\partial\mu}.$
Substituting from Eq. (12):
$\partial A_{\rho,\beta}(\mu)/\partial\rho=-\hat{v}(0)\sum_{n=1}^{\infty}nq_{n}e^{-n\rho\hat{v}(0)}e^{n\beta\mu},\quad\partial A_{\rho,\beta}(\mu)/\partial\mu=\beta\sum_{n=1}^{\infty}nq_{n}e^{-n\rho\hat{v}(0)}e^{n\beta\mu},\quad\mu=\partial f(\rho,\beta)/\partial\rho,$
we obtain
$\frac{\partial^{2}f(\rho,\beta)}{\partial\rho^{2}}=\frac{\lambda_{\beta}^{d}}{\beta\sum_{n=1}^{\infty}nq_{n}e^{-n\rho\hat{v}(0)}e^{n\beta\partial f(\rho,\beta)/\partial\rho}}+\frac{\hat{v}(0)}{\beta}.$ (17)
The interest in $\partial^{2}f(\rho,\beta)/\partial\rho^{2}$ lies in its relation with the $\rho$-derivative of the pressure:
@@
We now derive the $L\to\infty$ limit of $q^{L}_{n}$. As noted above, $q_{1}=1$ and $q_{2}=0$. For $n\geq 3$ we proceed as follows. Substituting $E_{L}(x)=1-v_{L}(x)$ in the formula for $\hat{E}_{L}(z)$,
@@
The dependence on $L^{d}$ shows that among $\{z^{j}_{i}\in{\mathbb{Z}}^{d}\setminus\{0\}|(ij)\in E(C)\}$ there are $N_{\rm free}(C)=|E(C)|-n+1$ free summation variables, and the residual n1n-1 variables are linear combinations of them. In the limit of $L$ going to infinity, $L^{-d(|E(C)|-n+1)}$ times the multiple sum over $\{z^{j}_{i}\in{\mathbb{Z}}^{d}\setminus\{0\}|(ij)\in E(C)\}$ becomes a multiple integral with respect to the free variables,
@@
This is in accordance with the blocks forming a tree,
@@
Remark 1. If a block is a planar graph or a graph consisting of the vertices and edges of a convex polyhedron, $N_{\rm free}$ is related to the Euler characteristic V+FE=2V+F-E=2. Comparison with it shows that $N_{\rm free}=F-1$, where $F$ is the number of faces (the exterior face included for planar graphs). We choose any F1F-1 faces of the polyhedron, or all the interior faces of the planar block, and assign a free variable to each bordering cycle. Referring to Eq. (19), $c_{ij}\neq 0$ for a single $j$ if $i$ is an external edge and for two $j$’s if $i$ is an internal edge.
Remark 2. For the complete $n$-graph the general solution of the system of equations $Z_{k}=0$, $k=1,\dots,n$, is easily obtained. In this case $N_{\rm free}=n(n-1)/2-n+1=(n-1)(n-2)/2$, the number of edges of the complete (n1)(n-1)-graph. The variables of this latter, $z^{k}_{j}$ for $1\leq j<k\leq n-1$ can be taken to be free; then, with
@@
all the equations will be satisfied. The integration variables are $\kappa^{k}_{j}$, $1\leq j<k\leq n-1$.
@@ 4 An illustrative example
$u$ depends on $\beta$, but as long as we only look at the $\rho$-dependence, the interaction is legitimate. Let us retain from $\Sigma_{n}$ only the (n1)!/2(n-1)!/2 cycles of length $n$. There are many more elements of $\Sigma_{n}$, and one can only hope that, with the cycles only, the result is qualitatively correct. There is no phase transition in one dimension for short-range repulsive interactions, so we focus on $d\geq 2$. For $n\geq 3$
@@
If $\beta\mu<d\ln(\lambda_{\beta}/\lambda)+\rho\hat{v}(0)$, the infinite sum is absolutely convergent. If $\beta\mu=d\ln(\lambda_{\beta}/\lambda)+\rho\hat{v}(0)$, it is still absolutely convergent for $d\geq 3$, nonabsolutely convergent for $d=1,2$, and for any $d\geq 1$ it is bounded below by $3^{-d/2}$, so
@@
is convergent, by Abel’s theorem $A_{\rho,\beta}(\mu)$ tends to it as $\mu$ tends to $\bar{\mu}(\rho,\beta)$. It remains for us to verify the consistency of the approximation. First, $A_{\rho,\beta}(\mu)$ must monotonically increase with $\mu$ for $\infty<\mu<\bar{\mu}(\rho,\beta)$; second, $\lim_{\mu\uparrow\bar{\mu}(\rho,\beta)}{\,\rm d}A_{\rho,\beta}(\mu)/{\,\rm d}\mu$ must exist, so that $\partial p(\rho,\beta)/\partial\rho$ has a left limit at $\rho_{c}(\beta)$, cf. Eqs. (17), (18). By introducing $\gamma=-\beta[\mu-\bar{\mu}(\rho,\beta)]$, $A_{\rho,\beta}(\mu)=(\lambda_{\beta}/\lambda)^{d}B(\gamma)$, where
@@ 5 No condensed phase without infinite clusters
Not knowing $q_{n}$ any better, we can nevertheless affirm that any condensed phase must contain infinite clusters. Consider, for example, the vapor-liquid transition. Fix a subcritical temperature, so $\beta>\beta^{*}$, $\bar{A}_{\beta}<\infty$. Increase the density of the vapor by increasing the pressure. At some $\rho=\rho_{v+}(\beta)$ condensation begins and, as we know from experiment, the van der Waals theory, and numerical simulations with, for example, the Lennard-Jones potential [15, 16], the vapor of density $\rho_{v+}(\beta)$ and the liquid of density $\rho_{l-}(\beta)$ coexist for the global density in the interval $(\rho_{v+}(\beta),\rho_{l-}(\beta))$. The chemical potential is the same in both phases and is therefore constant throughout the coexistence interval. It is $\mu(\rho_{v+}(\beta),\beta)=\epsilon(\beta)+\rho_{v+}(\beta)\hat{v}(0)/\beta$, cf. Eq. (15), where
@@
is valid for any $\rho\geq\rho_{v+}(\beta)$. Beyond some $\rho_{l-}(\beta)$ the chemical potential continues to increase with $\rho$, and the equation (27) defines a condition on it. By further increasing the pressure, we arrive at the coexistence of the liquid of density $\rho_{l+}(\beta)$ and a solid of density $\rho_{s-}(\beta)$. In the coexistence interval $(\rho_{l+}(\beta),\rho_{s-}(\beta))$ the chemical potential is again constant at the value $\partial f(\rho,\beta)/\partial\rho|_{\rho=\rho_{l+}(\beta)}$, and continues to increase beyond $\rho=\rho_{s-}(\beta)$. The infinite clusters in a solid must be different from those in the liquid. In a crystal the atomic positions are correlated on long distances, in the liquid the correlation is of medium range. This suggests that in a crystal the infinite clusters are macroscopic, while in the liquid they are submacroscopic. Accordingly, for $\rho\geq\rho_{v+}(\beta)$
@@
densities carried by finite, infinite submacroscopic, and macroscopic clusters. Submacroscopic clusters must be infinite in number to contribute to the density. If $\beta<\infty$ then $\rho_{v+}(\beta)>0$, finite clusters appear along with infinite clusters, in liquids and crystals, at positive temperatures. On the contrary, $\lim_{\beta\to\infty}\rho_{v+}(\beta)$ can be positive or zero, and both results have an interesting consequence on the ground state, the $\beta\to\infty$ limit of the infinite volume Gibbs state. If $\rho_{v+}(\infty)>0$, there are finite clusters in the ground state at any $\rho<\rho_{\rm cp}\leq\infty$, a possible result for short-range interactions. On the other hand, if $\rho_{v+}(\infty)=0$, there are only infinite clusters in the ground state at arbitrarily low density, a possible result for slowly decaying, but still integrable, interactions.